Quiz 4

\underline{Problem 1}: A simplified flowsheet for the Union Carbide oxo process is given below. The reactor operates at a pressure of 200 psia and a temperature of 373 K. The reaction mechanism is as follows:

P L + CO + H_{2} → IBA + NBA → HV

80% P L is converted and IBA/NBA ratio is 0.1. 1% of IBA is converted.

P L  Propylene  CO Carbon Monoxide
H_{2}  Hydrogen  IBA Isobutyraldehyde
NBA n-butyraldehyde  HV Heavy compounds
P  Propane

Propane is an inert component that does not take part in the reaction. The feed is available at 1 atm and 298 K and has the following composition:

CO 0.5 kgmol/s
H_{2} 0.5 kgmol/s
P L 0.47 kgmol/s
P  0.03 kgmol/s

Determine the overall conversion of propylene to n-butyraldehyde for a purge rate of 1%. Assume that all the separation steps (distillation) give perfect splits for the components shown.